blank1:

Jordan solved the equation −7x + 25 = 48; his work is shown below. Identify the error and where it was made. −7x + 25 = 48 Step 1: −7x + 25 − 25 = 48 − 25 Step 2: −7x = 23 Step 3: negative seven times x over seven = twenty-three over seven Step 4: x = twenty-three over seven

psirockin2:

In step 3, they divided -7x/7 and 23/7 They dropped their negative sign. so -x = 23/7 or x = -23/7

blank1:

thank you

psirockin2:

No problem